Skip to end of metadata
Go to start of metadata
  • tnw02: added last sonic; fixed cables-into-wrong-ports.  Now completely operational (#40).
  • rsw05: cycled power – hadn't been collecting data since 0145 this morning, even though had power.  Seems okay now.
  • tnw04: completely instrumented today.  Rewired INEGI's solar power system for 12VDC.  Now completely operational. (#41)
    • all sonics set with "box" facing the tower.
    • Top 2 levels (20m, 12m) set with boom to NW (to avoid hitting the solar panel while hauling up the sensor)

We're reasonably proud of ourselves!

 

 

 

  • No labels